1. Evaluate the following definite integral using the substitution formula: LI 4 cos(x) sin(x)dr.
Answer #1

27 COSK 4 cosx sinne da y= dy = Yet - sinu afe at The y = 1/2 at 25 y=1 I = S-14 ay [Fulting [putting the valuufy and dy 5 4y